OpenStudy (anonymous):

Which terms are considered like terms in the following expression? 3n+18 - 7n+3 + 5m can you please explain how you found the answer

OpenStudy (kainui):

"like terms" means they have a common factor. For instance suppose you have 3 rocks and 2 rocks. You can add these together to have 5 rocks, right? Similarly, if we just call our "rocks" as "x" instead you can see that we can collect these "like terms" as well! 3x+2x=5x just like 3 rocks + 2 rocks = 5 rocks. Now if you try to add these, don't be fooled, these are not like terms: x^2+x^3 can not add them because think about it, x squared and x cubed. You can't add squares with cubes, they're not the same!

OpenStudy (anonymous):

I'm still a little confused.

OpenStudy (kainui):

Sure, I'll help clear it up. So show me which terms you think are "like terms" and tell you what you need to do from there.

OpenStudy (anonymous):

3n and 7n im guessing.

OpenStudy (calculusfunctions):

@carlene Like terms are terms that have the same variables raised to the same exponents. To add or subtract them, you simply add or subtract their coefficients. e.g. -2x² and 4x². Their sum is 2x² (-2x² + 4x² = 2x²). Similarly 3x²y³ + x²y³ = 4x²y³. Unlike terms are terms that have different variables or the the same variables raised to different exponents. e.g. 3x and y or x²y³ and -x³y². You can't add or subtract these. Multiplying or dividing like or unlike terms is allowed, however you should familiarize yourself with the exponent laws.

OpenStudy (kainui):

Yep, so the like terms are: 3n, 7n, and 5n and 18 and 3 You can add these all up. So for example 3n+7n becomes 10n.

OpenStudy (calculusfunctions):

Yes 3n, -7n, and 5n would be like terms for example. You get it now. Similarly 18 and 3 are also like terms.

OpenStudy (anonymous):



OpenStudy (anonymous):

sorry that was suppose to be 5m

OpenStudy (calculusfunctions):

OH then only 3n and -7n are like terms, as well as 18 and 3. 5m will remain as 5m in the final answer but it won't be added to anything because there are no other m terms.

OpenStudy (anonymous):

Thanks everybody who helped i get it now.